# Probabilities of Chat LLMs Are Miscalibrated but Still Predict Correctness on Multiple-Choice Q&A

### v3

@@ 4 Calibration
Our first question was whether the MSPs of chat LLMs are calibrated. For each possible combination of LLM, dataset, and prompt phrasing, we performed two analyses. First, we computed each model’s calibration curve. For each model, we divided the range of possible MSPs into 10 quantile bins, i.e., each containing the same number of data points. Then for each bin, we computed the average MSP and the fraction of correct responses in that bin. Figure 2 displays each model’s calibration curve. Most models exhibit clear overconfidence: the MSP is consistently much larger than the fraction of correct responses. Most models are producing MSPs above 0.95 even when they are correct only 60% of the time.
Figure 2: The calibration curve for each model. Most models exhibit clear overconfidence: the MSP is larger than the true fraction of correct responses.
Figure 3: The data from Figure 2 is aggregated to a single data point per model: the model’s total calibration error (scaled by 100 to improve readability). There is no statistical evidence for a correlation between calibration error and Q&A accuracy ($r=-0.07,p=0.81$). The precise calibration error values can be found in the appendix (Table 3).
@@
Figure 3 shows that the answer is no. Although a slight downward trend exists, it is mostly due to three outlier models: Yi 34B and Llama 3.1 8B and 70B. [^8] Interestingly, the calibration error of the Llama 3.1 models is much lower than that of their 3.0 predecessors, despite the Q&A accuracy being very similar in the two families. More formally, the Pearson correlation coefficient (PCC) was $r=-0.07$ with $p=0.81$, indicating no correlation. [^9] PCCs and $p$ values for cross-model correlations (e.g, Figures 3 and 4) were computed via standard linear regression. As such, we should not expect chat LLMs to become more calibrated as their capabilities grow. Despite this, we will see in Section 5 that we can expect MSPs to become increasingly effective at predicting correctness, even as their calibration does not improve.
5 Predicting answer correctness with MSP and Max Logit
Before presenting those results, we must define correctness prediction. Correctness prediction is a binary classification task: given a multiple-choice question and the LLM’s response, predict whether the response is correct. We study two classifiers for this task: the MSP classifier (predict correctness iff the MSP exceeds some threshold) and the Max Logit [^10] Max Logit has no notion of calibration, since it is not a probability. classifier (predict correctness iff the maximum pre-softmax logit exceeds some threshold). We hypothesized that the MSP and Max Logit classifiers would perform (statistically) better than random chance on this classification task.
We are not training a new binary classifier for this task: we study the performance of the MSP and Max Logit “out of the box”, since our goal is to understand the innate properties of LLMs.
@@ 5.1 Methodology
We computed the AUROC for each possible combination of LLM, dataset, prompt phrasing, and classifier (MSP or Max Logit). We could not compute Max Logit AUROCs for the OpenAI models (GPT-3.5 Turbo and GPT-4o) because the OpenAI API only provides softmax probabilities and not pre-softmax logits. This resulted in 280 AUROC data points: 150 for MSP and 130 for Max Logit. For each model-classifier pair, we computed an unweighted average across 10 data points (five datasets $\times$ two prompt phrasings) to obtain the values in Table 1.
Statistical significance. To determine whether our AUROC results were statistically significant, we used the Mann-Whitney $U$ test (MWU) (Mann and Whitney, 1947; Wilcoxon, 1945). The MWU directly tests the null hypothesis that a classifier’s true AUROC is 50% (i.e., random guessing). For us, a significant MWU affirms the hypothesis that our classifier can distinguish between (1) questions where the LLM answered correctly and (2) questions where the LLM answered incorrectly.
For each available combination of model, dataset, prompt phrasing, and classifier (MSP and Max Logit), we tested the null hypothesis that the true AUROC was equal to 50%. This resulted in 280 Mann-Whitney $U$ tests. Table 1 reports the number of $p$-values which were below $10^{-4}$ for each model-classifier pair. The significance threshold of $\alpha=10^{-4}$ accounts for a Bonferroni correction (Bonferroni, 1936), which is applied when performing multiple hypothesis tests (in our case, 280) to ensure that the chance of falsely rejecting any null hypothesis is small. Starting from the standard threshold of $\alpha=0.05$, the Bonferroni correction yields $\alpha=0.05/280\approx 1.8\times 10^{-4}$. We use the stricter threshold of $10^{-4}$ for simplicity. [^12] We handle the cross-model comparisons separately since we test only 8 cross-model hypotheses (Figures 3 and 4, and 6 more in the appendix). As such, we use $\alpha=0.5/8\approx 0.006$ for those comparisons. This is not a crucial point since the $p$-values in Figures 3 and 4 are either above $0.3$ or below $10^{-3}$.
Table 1: Main AUROC results. AUROC and Q&A values are percentages, averaged over ten data points (five datasets and two phrasings). The $p<10^{-4}$ columns indicate how many of those ten data points yielded $p$-values below $10^{-4}$ for the null hypothesis that AUROC $=50\%$. The $p$-values are from the Mann-Whitney U test; see Section 5 for details.
@@
Falcon 7B 29.4 52.5 2/10 51.3 0/10
Falcon 40B 43.8 59.3 7/10 54.8 6/10
Llama 2 7B 40.4 56.9 6/10 55.7 5/10
Llama 2 70B 58.1 69.3 10/10 63.7 8/10
Llama 3.0 8B 58.5 71.3 10/10 68.9 10/10
Llama 3.0 70B 78.4 81.7 10/10 72.6 10/10
Llama 3.1 8B 62.8 72.7 10/10 67.1 10/10
Llama 3.1 70B 80.8 83.3 10/10 67.4 10/10
Mistral 7B 56.3 64.2 10/10 63.5 10/10
Mixtral 8x7B 69.0 60.2 9/10 61.2 10/10
SOLAR 10.7B 67.1 59.7 9/10 65.5 10/10
Yi 6B 53.0 66.6 10/10 60.5 10/10
Yi 34B 70.1 63.8 9/10 64.6 10/10
GPT-3.5 Turbo 67.5 75.6 10/10
GPT-4o 86.6 84.2 10/10
@@
Figure 4: Average AUROC vs average Q&A accuracy for the MSP (left) and Max Logit (right). These plots use the same data as Table 1. The PCCs for MSP and Max Logit were $r=0.81$ and $r=0.87$ respectively, both with $p<10^{-3}$, indicating strong correlations.
5.2 Key results
Among these 280 data points, AUROC outperformed random chance with $p<10^{-4}$ in 233 cases. When the Falcon and Llama 2 models are excluded, that statistic improves to 197/200. These results demonstrate that the MSP and Max Logit are statistically predictive of correctness (except for possibly the weakest models). However, this is perhaps not so surprising given the monotonicity of the calibration curves in Figure 2.
In our opinion, our more exciting finding is a strong direct correlation between average Q&A accuracy and average AUROC: the PCCs for MSP AUROC and Max Logit AUROC were $r=0.81$ and $r=0.87$, both with $p<10^{-3}$ (Figure 4). This finding suggests that as chat LLMs become more powerful, the uncertainty information present in MSPs actually becomes more refined. In contrast, we found no evidence that calibration error will similarly improve (Figure 3).
See Appendix A.2 for further discussion on our AUROC results. For example, AUROC values were much lower on WinoGrande, but Q&A accuracy was comparable to other datasets (Table 4).
@@ 6 Proof-of-concept: reducing wrong answers by abstention
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
[…]
Falcon 40B $-12.3$B1.21.21.21.2 $0.2$ $-68.5$ $-1.5$B0.5-0.5-0.5- 0.5
Llama 2 7B $-19.4$ $-11.1$B8.5-8.5-8.5- 8.5 $-79.0$ $-1.5$B0.1-0.1-0.1- 0.1
Llama 2 70B $16.2$B21.321.321.321.3 $19.1$ $-25.7$ $2.8$B3.03.03.03.0
Llama 3.0 8B $17.0$B23.523.523.523.5 $20.7$ $-24.4$B10.310.310.310.3 $10.0$
[…]
Llama 3.1 8B $25.6$B29.629.629.629.6 $25.3$ $-11.5$B18.018.018.018.0 $13.2$
Llama 3.1 70B $61.6$B62.862.862.862.8 $61.6$ $42.5$B52.052.052.052.0 $41.1$
Mistral 7B $12.7$B16.716.716.716.7 $15.1$ $-31.0$ $-5.0$B5.05.05.05.0
Mixtral 8x7B $38.1$B38.738.738.738.7 $34.2$ $7.2$ $12.7$B14.814.814.814.8
[…]
Yi 6B $6.0$B15.615.615.615.6 $9.7$ $-41.0$B6.56.56.56.5 $1.1$
Yi 34B $40.1$B41.141.141.141.1 $40.2$ $10.2$B14.214.214.214.2 $11.9$
GPT-3.5 Turbo $35.0$B38.038.038.038.0 $2.5$B27.927.927.927.9
GPT-4o $73.2$B73.573.573.573.5 $59.9$B61.461.461.461.4
Table 2: Results on Q&A with abstention. “Balanced” and “conservative” correspond to -1 and -2 points per wrong answer, respectively. Correct answers and abstentions are always worth +1 and 0 points, respectively. The total number of points is divided by the total number of questions (then scaled up by 100 for readability) to obtain the values shown in the table. We highlight the best method(s) for each model.
@@ 6.1 Methodology
These experiments use the same data as the AUROC experiments, but the data is analyzed differently. For each classifier (MSP or Max Logit) and a given threshold, we conducted the following analysis. First, we computed the classifier value (MSP or maximum logit) based on the initial LLM response, the same way we did in our AUROC experiments. If the classifier value was below the threshold, we recorded the model’s answer as “abstain” and otherwise recorded the original answer. We awarded 1 point per correct answer, 0 points per abstention, and $-c$ points per wrong answer, normalized by the total number of questions. We performed this analysis for $c=1$ (“balanced score") and $c=2$ (“conservative score"). For $c=1$, the benefit of a correct answer equals the cost of a wrong answer. However, wrong answers are often much worse (e.g., medical diagnoses), justifying $c=2$. Our experiment design was partly inspired by Kang et al. (2024), who use an even more extreme penalty of $c=4$.
Choosing the threshold. Unlike our AUROC results, here we must choose a specific threshold for whether the model should abstain. To do so, we randomly selected a training set of $k$ questions, and then for each model, chose the threshold which performed best on those $k$ questions. We discovered that $k=20$ performed almost as well as using half of all questions. Although $k=10$ still outperformed the base LLMs, performance was much worse than $k=20$. For this reason, we used $k=20$ in our main experiments. [^13] We also report results for $k=10$ (Table 6) and $k=$ half of all questions (Table 7) in the appendix. All tables and figures other than Tables 6 and 7 use $k=20$. Similar results can be obtained with similar values of $k$; the takeaway here is that a small amount of labeled data suffices for choosing an effective threshold.
@@ 6.2 Results
For each combination of LLM, classifier, and $c\in\{1,2\}$, Table 2 reports the scores obtained by the base LLM and by our method on the test set, where the our method used the threshold determined by training set. Figure 9 shows each model’s score across the entire range of possible thresholds.
@@
Our method outperformed or matched the base LLM in nearly all conditions and substantially outperformed the base LLM on the conservative score metric. (The sole exception is the Llama 3.1 8B, Max Logit, on the balanced score metric, which exhibits a tiny drop in score of 0.3). As expected, models with low initial scores exhibited the most dramatic improvements. For example, any model with a negative initial score can trivially improve to 0 by abstaining on every question. (Table 5 shows the abstention frequency for each model-classifier pair.) More generally, the higher the fraction of correct answers, the more likely the model is to accidentally abstain on a correct answer. As a result, it is unsurprising that models with high initial scores showed only modest improvements on the balanced score metric.
@@
These results show how the uncertainty signals from softmax probabilities and/or logits can be leveraged to improve performance on practical language tasks. Further details on our Q&A-with-abstention results can be found in Appendix A.3.

### v4

@@ 4 Calibration
We first asked whether the MSPs of chat LLMs are calibrated. For each possible combination of LLM, dataset, and prompt phrasing, we performed two analyses. First, we computed each model’s calibration curve as follows. For each model, we divided the range of possible MSPs into 10 quantile bins, i.e., each containing the same number of data points. Then for each bin, we computed the average MSP and the fraction of correct responses in that bin. Figure˜2 (left) displays each model’s calibration curve. Most models exhibit clear overconfidence: the MSP is consistently much larger than the fraction of correct responses. For example, most models produce MSPs above 0.95 even when they are correct only 60% of the time.
Figure 2: Left: The calibration curve for each model. Most models exhibit clear overconfidence: the MSP is larger than the true fraction of correct responses. Right: The average calibration error per model (scaled by 100 to improve readability) vs Q&A accuracy. There is no statistical evidence for a correlation between calibration error and Q&A accuracy ($R^{2}=0.07,p=0.33$). The precise calibration error values can be found in the appendix (Table˜6).
@@
Figure˜2 (right) shows that the answer is no. Although a slight downward trend exists, it is statistically insignificant. More formally, the coefficient of determination was $R^{2}=0.07$ ($p=0.33$), meaning that Q&A accuracy can explain only 7% of the variance in calibration error. [^8] Coefficients of determination and $p$ values for cross-model correlations (e.g., Figures 2 right and 3) were computed via standard linear regression. As such, we should not expect chat LLMs to become more calibrated as their capabilities grow. Despite this, we will see in Section˜5 that we can expect MSPs to become increasingly effective at predicting correctness, even as their calibration does not improve.
5 Correctness prediction without calibration
Before presenting said results, we must define correctness prediction. Correctness prediction is a binary classification task: given a multiple-choice question and the LLM’s response, predict whether the response is correct. We study two classifiers for this task: the MSP classifier (predict correctness iff the MSP exceeds some threshold) and the Max Logit [^9] Max Logit has no notion of calibration, since it is not a probability. classifier (predict correctness iff the maximum pre-softmax logit exceeds some threshold). We hypothesized that the MSP and Max Logit classifiers would (statistically) outperform random chance on this classification task.
Note that we are not training a new binary classifier for this task: we study the performance of the MSP and Max Logit “out of the box”, since our goal is to understand the innate properties of LLMs.
@@ 5.1 Methodology
We computed the AUROC for each available combination of LLM, dataset, prompt phrasing, and classifier (MSP or Max Logit). We could not compute Max Logit AUROCs for the OpenAI models (GPT-3.5 Turbo and GPT-4o) because the OpenAI API only provides softmax probabilities and not pre-softmax logits. This resulted in 280 AUROC data points: 150 for MSP and 130 for Max Logit. For each model-classifier pair, we computed an unweighted average across 10 data points (five datasets $\times$ two prompt phrasings) to obtain the values in Table˜2.
Statistical significance. To determine whether our AUROC results were statistically significant, we used the Mann-Whitney $U$ test (MWU) (Mann & Whitney, 1947; Wilcoxon, 1945). The MWU directly tests the null hypothesis that a classifier’s true AUROC is 50% (i.e., random guessing). For us, a significant MWU affirms the hypothesis that our classifier can distinguish between (1) questions where the LLM answered correctly and (2) questions where the LLM answered incorrectly.
For each available combination of model, dataset, prompt phrasing, and classifier, we tested the null hypothesis that the true AUROC was equal to 50%. This resulted in 280 MWUs. Table 2 reports the number of $p$-values which were below $10^{-4}$ for each model-classifier pair. The threshold of $\alpha=10^{-4}$ accounts for a Bonferroni correction (Bonferroni, 1936), which is applied when performing multiple hypothesis tests (in our case, 280) to ensure that the chance of falsely rejecting any null hypothesis is small. Starting from the standard threshold of $\alpha=0.05$, the Bonferroni correction yields $\alpha=0.05/280\approx 1.8\times 10^{-4}$. We use the stricter threshold of $10^{-4}$ for simplicity. [^11] We handle the cross-model comparisons separately since we test only 18 cross-model hypotheses (one in Figure 2, two in Figure 3, two in Figure 4, and 13 more in the appendix). As such, we use $\alpha=0.05/18\approx 0.003$ for those comparisons. For simplicity, we use $\alpha=10^{-3}$ for cross-model comparisons.
5.2 Results
Key results. Among these 280 data points, AUROC outperformed random chance with $p<10^{-4}$ in 245 cases. When the Falcon and Llama 2 models are excluded, that statistic improves to 196/200. These results demonstrate that the MSP and Max Logit are statistically predictive of correctness (except for possibly the weakest models).
Our more exciting finding is a strong direct correlation between average Q&A accuracy and average AUROC: the coefficients of determination for MSP AUROC and Max Logit AUROC were $R^{2}=0.94$ and $R^{2}=0.71$, both with $p<10^{-3}$ (Figure 3). This finding suggests that as chat LLMs become more powerful, the uncertainty information present in MSPs actually becomes more refined. In fact, that might be an understatement given that Q&A accuracy accounts for a remarkable 94% of the variance in MSP AUROC. In other words, a model’s Q&A accuracy almost entirely determines how well that model’s MSP can predict correctness, regardless of any other factors like architecture, size, etc. In contrast, we found no evidence that calibration error will similarly improve (Figure˜2, right).
Table 2: Main AUROC results. AUROC and Q&A values are percentages, averaged over ten data points (five datasets and two phrasings). The $p<10^{-4}$ columns indicate how many of those ten data points yielded $p$-values below $10^{-4}$ for the null hypothesis that AUROC $50\%$. The $p$-values are from the Mann-Whitney U test; see Section 5.1 for details.
@@ 5.2 Results
Falcon 7B 30.1 52.0 1/10 51.9 1/10
Falcon 40B 40.5 60.2 7/10 58.6 7/10
Llama 2 7B 39.7 58.6 6/10 57.8 7/10
Llama 2 70B 58.7 71.1 10/10 66.9 9/10
Llama 3.0 8B 59.1 71.2 10/10 68.9 10/10
Llama 3.0 70B 78.8 81.3 10/10 68.4 9/10
Llama 3.1 8B 62.0 72.9 10/10 67.3 10/10
Llama 3.1 70B 80.6 83.4 10/10 69.6 10/10
Mistral 7B 57.0 66.1 10/10 62.8 10/10
Mixtral 8x7B 69.9 70.0 10/10 60.5 8/10
SOLAR 10.7B 67.5 71.8 10/10 66.8 10/10
Yi 6B 49.6 67.0 10/10 60.4 10/10
Yi 34B 69.1 75.8 10/10 69.0 10/10
GPT-3.5 Turbo 67.5 76.1 10/10
GPT-4o 86.9 85.3 10/10
@@
Figure 3: Average AUROC vs average Q&A accuracy for the MSP (left) and Max Logit (right). These plots use the same data as Table 2. The coefficients of determination for MSP and Max Logit were $R^{2}=0.94$ and $R^{2}=0.71$ respectively, both with $p<10^{-3}$, indicating strong correlations.
One theory for the correlation between model capability and correctness prediction is the following:
  1. 1.
    For base models, calibration does improve with capability (Kadavath et al., 2022 and Section˜7).
  2. 2.
    Post-hoc calibration methods like Platt scaling (Platt, 2000) can improve calibration for miscalibrated LLMs (Shen et al., 2024; Ulmer et al., 2024; Xiao et al., 2025; Xie et al., 2024; Zhang et al., 2024).
  3. 3.
    Since AUROC depends only on the ordering of scores (in our case, the MSP or maximum logit) and Platt scaling preserves order, Points 1 and 2 suggest that correctness prediction AUROC will also improve with capability.
However, this line of reasoning does not explain the impressive strength of the correlation ($R^{2}=0.94$). Indeed, the correlation we find between Q&A accuracy and calibration error in base models in Figure˜7 is much weaker than this ($R^{2}=0.51,p=0.006$). We find this intriguing.
An outlier dataset: WinoGrande. WinoGrande was by far the hardest dataset for our correctness prediction task (Table 3). Our best hypothesis for this discrepancy is that WinoGrande is intentionally adversarial and tries to “trick” the model. An illustrative question from this dataset is “Neil told Craig that he has to take care of the child for the day because    did it last time." Even for some humans, it could be unclear whether Neil is assuming responsibility or assigning responsibility. One wrinkle is that the Q&A accuracy on WinoGrande is comparable to other datasets, so it is not the case that this dataset is “harder” in general: it is harder only for predicting correctness. Despite the average MSP AUROC of 60.8% for WinoGrande, Llama 3.1 70B and GPT-4o still achieved AUROCs of 75.8% and 78.3% respectively on this dataset (Table 25), suggesting that this difficulty is surmountable for capable models.
Minimal correlation between model size and AUROC. The coefficients of determination between model size and AUROC were $R^{2}=0.35$ ($p=0.03$) and $R^{2}=0.16$ ($p=0.17$) for MSP and Max Logit, respectively (Figure 4). It is unsurprising that some correlation exists, due to the known correlation between size and model capabilities (e.g., Q&A accuracy) and our correlation between Q&A accuracy and AUROC. However, the relatively weak correlation between model size and AUROC suggests that it is actually Q&A accuracy and not model size that matters.
Table 3: Average Q&A accuracy, AUROCs, and calibration error per dataset. All values are averaged over the 15 models and two prompts.
Q&A accuracy MSP AUROC Max Logit AUROC Calibration Error (x100)
ARC-Challenge 71.8 78.1 70.0 13.6
HellaSwag 60.1 71.0 63.9 19.6
MMLU 56.9 73.3 66.8 24.1
TruthfulQA 51.4 71.0 63.3 30.6
WinoGrande 60.8 60.8 54.8 24.2
Figure 4: AUROC vs model size for MSP (left) and Max Logit (right). The coefficients of determination for MSP and Max Logit were $R^{2}=0.35$ ($p=0.03$) and $R^{2}=0.16$ ($p=0.17$) respectively. GPT-3.5 Turbo and GPT-4o were excluded since their sizes are unknown.
Figure 5: Average AUROC vs Q&A accuracy based on prompt phrasing (the two phrasings can be found in Figures 1 and 8). All values are averaged over the five datasets.
Maximum vs margin. In the main body of the paper, we only consider the maximum probability and maximum logit. However, there are other ways to measure uncertainty as well. In Appendix˜C, we also consider the margin (difference between the top two probabilities/logits) and entropy (only for probabilities). One interesting finding is that for logits specifically, both the AUROC values and the strength of the correlation with Q&A accuracy are much higher for the margin compared to the maximum. This effect does not exist for probabilities. Our theory is that the probabilities are already normalized, while the logits are not, and the margin sort of functions as normalization.
Prompt phrasing had minimal impact. The two prompt phrasings (Figures 1 and 8) yielded similar results (Figure 5). This suggests that our results are robust to minor modifications to the prompt.
@@ 6.1 Methodology
These experiments use the same data as the AUROC experiments, but the data is analyzed differently. For each classifier (MSP or Max Logit) and a given threshold, we conducted the following analysis. First, we computed the classifier value (MSP or maximum logit) based on the initial LLM response, the same way we did in our AUROC experiments. If the classifier value was below the threshold, we recorded the model’s answer as “abstain” and otherwise recorded the original answer. We awarded 1 point per correct answer, 0 points per abstention, and c-c points per wrong answer, normalized by the total number of questions. We performed this analysis for $c=1$ (“balanced score”) and $c=2$ (“conservative score”). For $c=1$, the benefit of a correct answer equals the cost of a wrong answer. However, wrong answers are often much worse (e.g., medical diagnoses), justifying $c=2$. Our experiment design was partly inspired by Kang et al. (2024), who used an even more extreme penalty of $c=4$.
Choosing the threshold. Unlike our AUROC results, here we must choose a specific threshold for whether the model should abstain. To do so, we randomly selected a training set of $k$ questions per dataset. Then for each model, we chose the threshold which performed best on those $5k$ questions. We use a single threshold for each model across all datasets in order to make our method more robust. We discovered that $k=20$ performed almost as well as using half of all questions. Figure˜10 in the appendix visualizes the Q&A-with-abstention score as a function of $k$. Unless otherwise specified, all figures and tables use $k=20$.
@@ 6.2 Results
Table 4: Results on Q&A with abstention. “Balanced” and “conservative” correspond to -1 and -2 points per wrong answer, respectively. Correct answers and abstentions are always worth +1 and 0 points, respectively. The total number of points is divided by the total number of questions (then scaled up by 100 for readability) to obtain the values shown in the table. We highlight the best method for each model.
Balanced Conservative
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B $39.8 \B$0.7-0.7 2.2-2.2 $109.7 \B$7.8-7.8 8.0-8.0
[…]
Yi 6B $0.9 \B$$9.6$ $5.0$ $51.4 \B$$5.8$ 0.2-0.2
[…]
@@
For each combination of LLM, classifier, and $c\in\{1,2\}$, Table 4 reports the scores obtained by the base LLM and by our method on the test set, where our method used the threshold determined by the training set. Figure 9 shows each model’s score across the entire range of possible thresholds. Our method outperformed or matched the base LLM in all conditions and substantially outperformed the base LLM on the conservative score metric.
@@
As expected, models with low initial scores exhibited the most dramatic improvements. For example, any model with a negative initial score can trivially improve to 0 by abstaining on every question. More generally, the higher the fraction of correct answers, the more likely the model is to accidentally abstain on a correct answer. As a result, it is unsurprising that models with high initial scores showed more modest improvements and lower abstention rates.
Abstention frequency. In line with the reasoning above, some models do abstain quite frequently, with some of the weakest models reaching nearly 100% abstention rates (Table 7 in the appendix). One may be concerned that excessively frequent abstention could render a model unusable. However, we argue that excessively frequent wrong answers would render a system not only unusable but actively harmful. If a model is likely to be wrong more often than not, perhaps it is appropriate for the model to always abstain.
@@
Overall, our Q&A-with-abstention results show how the uncertainty signals from softmax probabilities and/or logits can be leveraged to improve performance on practical language tasks.
7 Experiments on base LLMs
Our paper focuses on chat LLMs since the questions we study are relatively better understood for base LLMs (due to Kadavath et al., 2022 in particular). However, given our goal of providing a comprehensive study of correctness-awareness, we ran the same suite of experiments and same analysis (calibration, AUROC, and Q&A-with-abstention) for the base LLMs. This section overviews our results for the base models, with details appearing in Section˜A.3. The OpenAI models are excluded from these experiments since we do not have access to their corresponding base models.
The base models were significantly worse at responding in the correct format. That is, it was not uncommon for a model to assign very small probabilities to all of the possible answer tokens (Table˜5). The answer token probabilities were normalized to 1 before analysis (see Section˜3), but we think the base models’ results should be regarded with mild skepticism. That said, our findings do align with Kadavath et al. (2022).
Table 5: Median unnormalized MSPs for base LLMs.
LLM Median unnormalized MSP
Falcon 7B 0.002
Falcon 40B 0.023
Llama 2 7B 0.0037
Llama 2 70B 0.020
Llama 3.0 8B 0.099
Llama 3.0 70B 0.102
Llama 3.1 8B 0.124
Llama 3.1 70B 0.078
Mistral 7B 0.285
Mixtral 8x7B 0.187
SOLAR 10.7B 0.020
Yi 6B 0.182
Yi 34B 0.005
Figure 6: Calibration results for the base models. The base models are much better calibrated than the chat models, and calibration of base models does improve as Q&A accuracy improves ($R^{2}=0.51,p=0.006$).
Figure 7: AUROC results for the base models for MSP (left) and Max Logit (right). We see strong correlations between Q&A accuracy and both MSP AUROC ($R^{2}=0.87$) and Max Logit AUROC ($R^{2}=0.68$), both with $p<10^{-3}$.
Calibration. Figure˜6 shows that base models are fairly well calibrated and that calibration further improves with Q&A accuracy ($R^{2}=0.51,p=0.006$). Note that $p=0.006$ does exceed our Bonferroni-adjusted significance threshold of $10^{-3}$. However, we think this is less of a concern for this result since these findings mirror those of Kadavath et al. (2022). However, it is notable that this correlation is much weaker than the correlations between AUROC and Q&A accuracy, both for chat models (Figure˜3) and base models (Figure˜7).
AUROC. Figure˜7 shows the AUROC results for base models, analogous to Figure˜3. We see the same strong correlation between Q&A accuracy and both MSP AUROC ($R^{2}=0.87$) and Max Logit AUROC ($R^{2}=0.68$), both with $p<10^{-3}$. We observe that the correlations are not quite as strong as for the chat models (Figure˜3), and the AUROC values themselves are somewhat lower, although this may be noise. We also note that the base models’ correlations between Q&A accuracy and AUROC are much stronger than the base models’ correlation between Q&A accuracy and calibration error.